A cannonball on the ground is fired at 35° with an initial velocity of 250 m/s.


a. How long will it take to hit the ground?

b. How far from the cannonball will hit the ground?

c. Compute for the maximum height to be reached by the cannonball.



1
Expert's answer
2021-04-26T16:59:29-0400

a. How long will it take to hit the ground?


"t=2\frac{v_y}{g}=\frac{2v\sin\theta}{g}=29.3\text {s}."

b. How far from the cannonball will hit the ground?


"R=v_xt=v\cos\theta·t=5993\text{ m}."

c. Compute for the maximum height to be reached by the cannonball.


"h_\text{max}=\frac{v_y^2}{2g}=\frac{v^2\sin^2\theta}{2g}=1049\text{ m}."